OpenStudy (anonymous):

3(m + 5 + 9m) I have to find two equivalent expressions for this expression can someone tell me if there correct?

OpenStudy (anonymous):

i already have one its 15+30m

OpenStudy (zzr0ck3r):

\(3(m+5+9m)=3m+15+27m = 15+3m+27m=15+m(3+27)=\\3(5+m(1+9))\)